OpenStudy (anonymous):

10x-10-7x is greater than or equal to 3x-2

OpenStudy (anonymous):

first of all simplify : \(10x-10-7x\) we can write : 10x-10-7x as 10x-7x-10 .

OpenStudy (anonymous):

Can you tell me that what is 10x-7x-10 ?

OpenStudy (anonymous):

It would be 3x-10

OpenStudy (anonymous):

excellent! So now, we have : \(3x-10 \ge 3x-2 \) Correct?

OpenStudy (anonymous):

Yeah I got that far

OpenStudy (anonymous):

so now, 3x and 3x will be canceled out, do you know why?.

OpenStudy (anonymous):

Both would be a positive to each side

OpenStudy (anonymous):

Sorry negative

OpenStudy (anonymous):

like this : \(3x - 10 \ge 3x - 2\) \(\implies 3x-10-3x \ge 3x-3x-2 = -10 \ge -2 \) What I did was : Just subtracted 3x both sides.

OpenStudy (anonymous):

OK?

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

so can you tell me, is -10 equal or greater to -2 ?

OpenStudy (anonymous):

Hey, wait! can I have exact question please, are any options given?

OpenStudy (anonymous):

-10 is less than or equal to -2

OpenStudy (anonymous):

Answers are X >= -8 X<= 8 All real numbers No solution

OpenStudy (anonymous):

this first one is 3x-10, so 3x-10<3x-2, for example, then x=0,we can get -10<-2 is true right?